OpenStudy (anonymous):

Factor completely: 3x^2 − x − 4 A.(3x – 1)(x + 4) B.(3x + 4)(x – 1) C.(3x – 2)(x + 2) D.(3x – 4)(x + 1)

OpenStudy (anonymous):

3x^2 - x - 4 (3x- 4)(x + 1) check using the foil method 3x times x = 3x^2 3x times 1 = 3x -4 times x = -4x -4 times 1 = -4 now put it together... 3x^2 + 3x - 4x - 4 (combine like terms) 3x^2 - x - 4 (correct) D is your answer